OpenStudy (anonymous):

25xto the 3rd degreey to 2degree +125xto 2 degreeyto5degree FACTOR COMPLETELY

OpenStudy (anonymous):

\[(25x^3)^2+(125x^2)^5\]?

OpenStudy (anonymous):

HOW DO YOU ENTER EXPONENTS ON THE KEYBOARD

OpenStudy (anonymous):

i think you have a symol palette where it says \[\sum\]equation

OpenStudy (anonymous):

i am writing latex to do it

OpenStudy (anonymous):

i would start by writing \[(5^2x^3)^2+(5^3x^2)^5\]

OpenStudy (anonymous):

if this is the problem.

OpenStudy (anonymous):

if it is not let me know and i will not continue with this method

OpenStudy (anonymous):

NO WHEN I PUT TO WHATEVER DEGREE THAT IS THE EXPONENT

OpenStudy (anonymous):

25XTO 3 DEGREE Y TO 2 DEGREE + 125X TO 2 DEGREE Y TO 5 DEGREE

OpenStudy (anonymous):

\[25(x^3)^2+125(x^2)^5\]?

OpenStudy (anonymous):

YOU LEFT OUT HE Y'S AND THEIR DEGREES AND NO BRACKETS OR PARENTHESES

OpenStudy (anonymous):

oho

OpenStudy (anonymous):

\[25x^3y^2+125x^2y^5\]

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

ok each has a common factor of \[25x^2y^2\] is that clear?

OpenStudy (anonymous):

so you can factor it out and write \[25x^2y^2(x+5y^3\]

OpenStudy (anonymous):

NO, I NEED STEP BY STEP

OpenStudy (anonymous):

ok. you have two numbers 25 and 125

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

their greatest common factor is 25 because \[25=25\times 1\] and \[125=25\times 5\] so that is going to come out of the parenthese

OpenStudy (anonymous):

just looking at that part we can say that \[25+125=25(1+5)\]

OpenStudy (anonymous):

THERE IS NO PARENTHESES

OpenStudy (anonymous):

no there are not. but "factoring" means to write as a produce. that is our job, to put the parenthese in

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

so now for the variables: first term has \[x^3\] second term has \[x^2\]

OpenStudy (anonymous):

their greatest common factor is \[x^2\] because \[x^2=x^2\times 1\] and \[x^3=x^2\times x\]

OpenStudy (anonymous):

I AM SORRY BUT THIS ISN'T HELPING

OpenStudy (anonymous):

so we are going to "factor out" a 25 and an \[x^2\]

OpenStudy (anonymous):

i have a better idea.

OpenStudy (anonymous):

look at my answer which is \[25x^2y^2(x+5y^3)\] multiply out using the distributive law and see if you get what you started with.

OpenStudy (anonymous):

maybe then it will be clear where the \[25x^2y^2\] came from, and why we pulled it out front of the parentheses